# Illuminating Black Hole Shadow with Dark Matter Annihilation

### v2

@@ IV Discussion
The exceptional angular resolution offered by the EHT and the ngEHT not only enriches our understanding of astrophysical phenomena but also provides robust constraints on fundamental particle physics Chen et al. (2020); Yuan et al. (2021); Chen et al. (2022a, b, 2023); Ayzenberg et al. (2023). This study illustrates how the detailed morphology of BH images, particularly the inner shadows captured by the ngEHT, can inform constraints on DM annihilation. We capitalize on the heightened DM density near SMBHs and the unique characteristics of BH images—most notably, the diminished intensity in the inner shadow area, a result of the jet’s sparse electron density. To ensure the reliability of our constraints, we have developed an advanced framework for simulating the propagation of electrons and positrons within a background fit derived from a realistic GRMHD profile. This profile is consistent with the latest intensity and polarimetric observations and represents a advancement beyond previously assumed simpler spherical accretion models Regis and Ullio (2008); Lacroix et al. (2014, 2017); Yuan et al. (2022). While our analysis uses M87 as a primary example, our discussions and constraints are equally applicable to another EHT focus, Sgr A, within the Milky Way—given that observations similarly support a MAD profile around it and an almost direct line of sight Akiyama et al. (2022); Abuter et al. (2023).
@@
Looking ahead, this research aims to extensively explore DM characteristics, including an expanded range of annihilation channels, a wider spectrum of mass scales, and exploring phenomena such as p-wave annihilation Christy et al. (2023) and forbidden DM Cheng et al. (2023), both of which are expected to be significantly amplified in the vicinity of SMBHs. Furthermore, our analysis of the horizon-scale intensity map could be broadened to assess morphological changes at larger scales, encompassing extended jet structures and the accretion flow at greater distances from the BH. In these areas, the electron density from DM annihilation with a large cross-section decreases more slowly than the astrophysical background.
@@
From an observational standpoint, the incorporation of linear and circular polarization intensity measurements could unveil new perspectives in differentiating between electrons and positrons Anantua et al. (2020); Emami et al. (2021, 2023); Anantua et al. (2024). This differentiation is crucial since astrophysical plasmas typically exhibit a significant deficit in positron populations, potentially facilitating the imposition of tighter constraints. The advent of forthcoming multi-frequency observations Chael et al. (2023) is expected to further improve detection capabilities.
Acknowledgements.
@@
We are grateful to Richard Anantua, Marco Cirelli, Razieh Emami, Jordan Koechler, George Wong, Xiao Xue, and Yue Zhao for useful discussions. Y.C. is supported by VILLUM FONDEN (grant no. 37766), by the Danish Research Foundation, and under the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587, and by the European Consortium for Astroparticle Theory in the form of an Exchange Travel Grant, and by FCT (Fundação para a Ciência e Tecnologia I.P, Portugal) under project No. 2022.01324.PTDC. R.D. is supported in part by the National Key R&D Program of China (No. 2021YFC2203100). Y.M. is supported by the National Natural Science Foundation of China (Grant No. 12273022), the Shanghai Municipality orientation program of Basic Research for International Scientists (grant no. 22JC1410600), and the National Key R&D Program of China (No. 2023YFE0101200). J.S. is supported by Peking University under startup Grant No. 7101302974 and the National Natural Science Foundation of China under Grants No. 12025507, No.12150015; and is supported by the Key Research Program of Frontier Science of the Chinese Academy of Sciences (CAS) under Grants No. ZDBS-LY-7003. Y.C. and J.S. are supported by the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP), which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy – EXC-2094 – 390783311.
@@
The supplemental materials provide a comprehensive procedure for calculating the electron-positron spectrum outside the black hole (BH) resulting from dark matter (DM) annihilation, as well as the intensity map derived from synchrotron radiation emitted by these electrons and positrons. Additionally, the formalism for the critical curve and the contour of the inner shadow is included. Throughout this study, we employ natural units.
@@ I.1 Propagation of Electrons and Positrons
The initial terms on the left-hand side represent the advective transport of the electron-positron flow and energy acquisition through adiabatic compression, indicated by the rate of momentum change $\dot{\vec{p}}_{\rm adi}$, where $\vec{v}_{b}$ is the flow’s bulk velocity. The subsequent term addresses energy depletion via assorted radiative processes, marked by the rate of momentum change $\dot{\vec{p}}_{\rm rad}$. The concluding terms introduce spatial diffusion and diffusive reacceleration impacts, with $D_{xx}$ and $D_{pp}$ symbolizing the diffusion coefficients in spatial and momentum spaces, respectively. The term $q(\vec{r},p)$ signifies the DM source function for electrons and positrons.
Considering the relativistic speeds of electrons and positrons far surpass the flow’s bulk velocity $\vec{v}_{b}$, and that the DM source function $q(\vec{r},p)$ is isotropic in momentum space—as are the momentum change processes $\dot{\vec{p}}_{\rm adi}$ and $\dot{\vec{p}}_{\rm rad}$, which show no directional preference under statistical mean—it is plausible to assume that $f_{e}$ exhibits isotropy in momentum space. This leads to a simplification where the dependency on $\vec{p}$ narrows down to reliance on $p$ alone.
The rate of momentum gain from adiabatic compression is given by $\dot{p}_{\rm adi}=-p(\nabla\cdot\vec{v}_{b}\,)/3$. Energy losses stem from various processes, including synchrotron radiation in the presence of a background magnetic field, bremsstrahlung scattering with background ions, inverse Compton scattering with background radiation, and Coulomb scattering with background electrons. These losses are collectively denoted by $\dot{p}_{\rm rad}=\dot{p}_{\rm syn}+\dot{p}_{\rm brem}+\dot{p}_{\rm IC}+\dot{p}_{\rm C}$. Assuming typical astrophysical conditions such as a magnetic field strength of $B\sim 1\,\rm G$, a number density of background ions and thermal electrons of $\bar{n}_{ion/e}\sim 1\,\rm cm^{-3}$, and an energy density of background radiation of $\bar{u}_{\rm r}\sim 1\,\rm eV\,cm^{-3}$, the estimated orders of magnitude for various rates of momentum change are Regis and Ullio (2008); Chael et al. (2017):
$\begin{split}\dot{p}_{\rm adi}&\approx 1.03\times 10^{-6}\,{\rm GeV\,s^{-1}}\left(\frac{p}{1\mathrm{GeV}}\right)\left(\frac{r}{r_{g}}\right)^{-1}\left(\frac{v_{b}}{0.1}\right)\,,\\ \dot{p}_{\rm syn}&\approx-2.01\times 10^{-7}\,{\rm GeV\,s^{-1}}\left(\frac{B}{1\,\rm G}\right)^{2}\left(\frac{p}{1\,\rm GeV}\right)^{2}\,,\\ \dot{p}_{\rm brem}&\approx-1.37\times 10^{-16}\,{\rm GeV\,s^{-1}}\left(\frac{\bar{n}_{ion}}{1\,\rm cm^{-3}}\right)\left(\frac{p}{1\,\rm GeV}\right)\left(\ln\frac{p}{1\,\rm GeV}+7.94\right)\,,\\ \dot{p}_{\rm IC}&\approx-1.02\times 10^{-16}\,{\rm GeV\,s^{-1}}\left(\frac{\bar{u}_{\rm r}}{1\,\rm eV\,\rm cm^{-3}}\right)\left(\frac{p}{1\,\rm GeV}\right)^{2}\,,\\ \dot{p}_{\rm C}&\approx-7.62\times 10^{-18}\,{\rm GeV\,s^{-1}}\left(\frac{\bar{n}_{e}}{1\,\rm cm^{-3}}\right)\left[\ln\frac{p}{1\,\rm GeV}+\ln\frac{\bar{n}_{e}}{1\,\rm cm^{-3}}+82.3\right]\,,\end{split}$ (S2)
where we take the reference length to be the gravitational radius of M87, $r_{g}=3.1\times 10^{-4}$ pc. The signs ++/-+ / - represent momentum gain or loss, respectively. For adiabatic compression, we have approximated $-\nabla\cdot\vec{v}_{b}\approx v_{b}/r$. It emerges that $\dot{p}_{\rm adi}$ and $\dot{p}_{\rm syn}$ provide the most substantial contributions, while the effects of the other terms can be considered negligible in Eq. (S1).
The spatial diffusion term is quantified by the distance over which charged particles dissipate most of their energy while diffusing, denoted as $d_{L}\equiv(D_{xx}p/\dot{p})^{1/2}$ Regis and Ullio (2008). In the context of Bohm diffusion, where the diffusion coefficient for electrons is given by $D_{xx}=p/(3eB)$, and considering energy losses primarily due to synchrotron radiation ($\dot{p}=\dot{p}_{\rm syn}$), the characteristic diffusion length $d_{L}$ can be approximated as
@@
$\displaystyle d_{L}\approx 4.18\times 10^{-4}\,r_{g}\,\left(\frac{B}{1\mathrm{G}}\right)^{-3/2},$ (S3)
which is significantly less than the typical propagation lengths of electrons and positrons, estimated to be on the order of $r_{g}$. In the context of Eq. (S1), the comparative significance of the spatial diffusion relative to be synchrotron radiation is approximated by the ratio $d_{L}/r_{g}$. Therefore, for the purpose of our analysis, the impact of spatial diffusion can be overlooked.
The diffusive reacceleration term originates from magnetic turbulence, where stochastic acceleration by Alfvén waves contributes to the additional energy gain of charged particles. Assuming the magnetic turbulence follows a power-law spectrum with index $\alpha$, the momentum-diffusion coefficient behaves as $D_{pp}\sim\beta_{A}^{2}\zeta r_{L}^{\alpha-2}\lambda_{\rm max}^{1-\alpha}p^{\alpha}$ Dermer et al. (1996). Here, $\beta_{A}\approx 1$ represents the Alfvén velocity, $\zeta\equiv\delta B^{2}/B^{2}$ denotes the ratio of turbulent to average magnetic field strength, $r_{L}\equiv m/(eB)$, and $\lambda_{\rm max}$ is the maximum wavelength of the Alfvén waves. The rate of momentum gain associated with this process is expressed as
@@
$\displaystyle\dot{p}_{\rm acc}=\frac{m_{e}}{p^{2}}\frac{\partial}{\partial p}\left(p^{2}D_{pp}\right)\sim m_{e}\beta_{A}^{2}\zeta\,r_{L}^{\alpha-2}\lambda_{\rm max}^{1-\alpha}p^{\alpha-1}\approx 1.52\times 10^{-9}\,{\rm GeV\,s^{-1}}\beta_{A}^{2}\left(\frac{\zeta}{0.1}\right)\left(\frac{10^{4}\,r_{g}}{\lambda_{\rm max}}\right)\left(\frac{p}{\mathrm{GeV}}\right)\,,$ (S4)
adopting benchmark parameters of $\alpha=2$, $\zeta\sim 0.1$, and $\lambda_{\rm max}\sim 10^{4}\,r_{g}$. The contribution of this term, when compared to $\dot{p}_{\rm adi}$ and $\dot{p}_{\rm syn}$ within the framework of Eq. (S1), is significantly smaller and thus can be disregarded in our analysis.
@@
Synthesizing the discussion above, we retain only the first three terms in Eq. (S1) and simplify the $\vec{p}$-dependence for the isotropic distribution. Consequently, Eq. (S1) can be reformulated as
@@
which characterizes the evolution of the electron-positron flow along specific streamlines, defined by the bulk velocity $\vec{v}_{b}$. Given their substantial charge-to-mass ratios, electrons and positrons closely follow the magnetic field lines. As a result, at any given point along their path, the streamline direction $\hat{v}_{b}(\vec{r}\,)$ aligns with the magnetic field direction, denoted as $\hat{B}(\vec{r}\,)$.
For convenience, we define $s$ as the distance traveled along a streamline, parameterized by $\vec{r}\,(s)$. This notation helps express the rate of momentum change for electrons and positrons during their journey:
@@
which can be integrated to obtain:
$\displaystyle p-p_{\rm inj}(s_{\rm inj})=\int_{s_{\rm inj}}^{s_{\vec{r}}}{\rm d}s\,\frac{\dot{p}_{\rm adi}+\dot{p}_{\rm syn}}{v_{b}}.$ (S7)
Here, $s_{\vec{r}}$ represents the location at $\vec{r}$. $s_{\rm inj}$ is a parameter moving along the streamline, marking the starting points of a specific flow in phase space originating from DM injections, to be discussed in the next subsection, and $p_{\rm inj}$ is the corresponding momentum of the injection.
To resolve Eq. (S5), we first express its left-hand side as ${\rm d}f_{e}/{\rm d}s$, and then aggregate contributions from all possible flows propagating to $\vec{r}$:
$\displaystyle f_{e}(\vec{r},p)=\int^{s_{0}}_{s_{\vec{r}}}{\rm d}s_{\rm inj}\,\frac{q\,(s_{\rm inj},p_{\rm inj}(s_{\rm inj}))}{v_{b}(s_{\rm inj})}\exp\left[\int^{s_{\rm inj}}_{s_{\vec{r}}}{\rm d}{s}\,\frac{4\,\dot{p}_{\rm syn}(s,p_{\rm inj}({s}))}{p_{\rm inj}({s})\,v_{b}({s})}\right]\,,$ (S8)
where $s_{0}$ signifies the most distant injection point along a given streamline. The relationship between $p_{\rm inj}$ and $s$ is determined by Eq. (S7). By incorporating Eq. (S6) into Eq. (S8), the phase space distribution function can be articulated as:
$\displaystyle f_{e}(\vec{r},p)=\int^{s_{0}}_{s_{\vec{r}}}{\rm d}s_{\rm inj}\frac{q\,(s_{\rm inj},p_{\rm inj}(s_{\rm inj}))}{v_{b}(s_{\rm inj})}\left(\frac{p_{\rm inj}(s_{\rm inj})}{p}\right)^{4}G(s_{\vec{r}})\,G^{-1}(s_{\rm inj}),$ (S9)
where $G$ denotes a form factor satisfying:
@@
The term $G(s_{\vec{r}})\,G^{-1}(s_{\rm inj})$ in Eq. (S9) delineates the contraction or dilation of the volume in the configuration space as particles are translocated from their injection point $\vec{r}_{\rm inj}(s)$ to $\vec{r}$ along the streamline. As delineated by Eq. (S11), $G$ encompasses two distinct elements: the first stems from the convergence or divergence of $\hat{v}_{b}$, which aligns with the magnetic field lines; the second represents how variations in bulk velocity influence particle density. In a case of spherically symmetric magnetic fields and bulk velocities akin to free-fall, it is posited that $G(s_{\vec{r}})\propto r^{-2}$, thereby corroborating the results in Refs. Aloisio et al. (2004); Regis and Ullio (2008).
For the sake of simplicity, this study does not incorporate general relativistic effects within the transport equation, an approach that remains valid for larger radii where $r\gg r_{g}$. Given that the significant contributions to $f_{e}(\vec{r},p)$ primarily originate from injections traversing the outer region—even for locations $\vec{r}$ in proximity to the event horizon—this assumption has a minimal impact on the numerical results.

### v3

@@ Abstract
Discussion–The exceptional angular resolution offered by the EHT and its future upgrades not only enriches our understanding of astrophysical phenomena but also provides robust constraints on fundamental particle physics Chen et al. (2020); Yuan et al. (2021); Chen et al. (2022a, b, 2023); Ayzenberg et al. (2025). This study illustrates how the detailed morphology of BH images, particularly the shadow regions, can inform constraints on DM annihilation. We capitalize on the heightened DM density near SMBHs and the unique characteristics of BH images–most notably, the diminished intensity in the inner shadow area, a result of the jet’s sparse electron density. To ensure the reliability of our constraints, we have developed an advanced framework for simulating the propagation of electrons and positrons within a background fit derived from a realistic GRMHD profile. This profile is consistent with the latest intensity and polarimetric observations and represents an advancement beyond previously assumed simpler spherical accretion models Regis and Ullio (2008); Lacroix et al. (2014, 2017); Yuan et al. (2022). While our analysis uses M87 as a primary example, our discussions and constraints are equally applicable to another EHT focus, Sgr A, within the Milky Way–given that observations similarly support a MAD profile around it and an almost direct line of sight Akiyama et al. (2022); Abuter et al. (2023). Note that the DM considered here contributes solely to intensity enhancement, with negligible effects on geodesics due to its minimal mass distribution near the horizon. This distinguishes it from other dark sector phenomena, such as superradiant clouds Sengo et al. (2023); Chen et al. (2023), which can reach up to $10\%$ of the BH’s mass, or BH alternatives Sengo et al. (2024); Huang et al. (2024).
@@
Looking ahead, this research aims to extensively explore DM characteristics, including an expanded range of annihilation channels, a wider spectrum of mass scales, and exploring phenomena such as $p$-wave annihilation Christy et al. (2023) and forbidden DM Cheng et al. (2023), both of which are expected to be significantly amplified in the vicinity of SMBHs. Furthermore, our analysis of the horizon-scale intensity map could be broadened to assess morphological changes at larger scales, encompassing extended jet structures and the accretion flow at greater distances from the BH. In these areas, the electron density from DM annihilation with a large cross section decreases more slowly than the astrophysical background.
@@
From an observational standpoint, the incorporation of linear and circular polarization intensity measurements could unveil new perspectives in differentiating between electrons and positrons Anantua et al. (2020); Emami et al. (2021, 2023); Anantua et al. (2024). This differentiation is crucial since astrophysical plasmas typically exhibit a significant deficit in positron populations, potentially facilitating the imposition of tighter constraints. The advent of forthcoming multifrequency observations Chael et al. (2023) is expected to further improve detection capabilities.
@@
Acknowledgments–We are grateful to Richard Anantua, Marco Cirelli, Razieh Emami, Jordan Koechler, George Wong, Xiao Xue, Daixu Yang and Yue Zhao for useful discussions. The Center of Gravity is a Center of Excellence funded by the Danish National Research Foundation under Grant No. 184. Y. C. acknowledges support by VILLUM Foundation (Grant No. VIL37766) and the DNRF Chair program (Grant No. DNRF162) by the Danish National Research Foundation, the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” Grant Agreement No. Gravitas–101052587, the Rosenfeld foundation in the form of an Exchange Travel Grant, and the COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology). This project has received funding from the European Union’s Horizon 2020 research and innovation programm under the Marie Sklodowska-Curie Grant Agreement No. 101007855 and Grant Agreement No. 101131233. R.D. is supported in part by the National Key R&D Program of China (No. 2021YFC2203100). Y.M. is supported by the National Natural Science Foundation of China (Grant No. 12273022), the Shanghai Municipality orientation program of Basic Research for International Scientists (Grant No. 22JC1410600), and the National Key R&D Program of China (No. 2023YFE0101200). J.S. is supported by the National Key Research and Development Program of China under Grants No. 2020YFC2201501 and No. 2021YFC2203004, Peking University under startup Grant No. 7101302974, and the NSFC under Grants No. 12025507, No. 12150015, and No. 12450006.
Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
@@
The Supplemental Material provides a comprehensive procedure for calculating the electron-positron spectrum outside the black hole (BH) resulting from dark matter (DM) annihilation, as well as the intensity map derived from synchrotron radiation emitted by these electrons and positrons. Additionally, the formalism for the critical curve and the contour of the inner shadow is included. We also demonstrate that astrophysical uncertainties have only a minor impact on our results. Throughout this study, we employ natural units.
@@ I.1 Propagation of Electrons and Positrons
Here, $\vec{v}_{b}$ represents the bulk flow velocity of the plasma, while $\dot{p}_{\rm adi}$ and $\dot{p}_{\rm rad}$ denote the momentum change rates due to adiabatic compression/expansion and radiation, respectively. The terms $D_{xx}$ and $D_{pp}$ correspond to the spatial and momentum diffusion coefficients, and $q(\vec{r},p)$ is the source term for DM annihilation. The first two terms describe advective transport and adiabatic compression, while the remaining terms account for radiative losses and diffusion.
We can simplify Eq. (S1) using the isotropic momentum approximation, as the relativistic electrons and positrons far exceed the bulk flow velocity $\vec{v}_{b}$, and both the DM source function $q(\vec{r},p)$ and the relevant momentum change processes are isotropic in momentum space.
Below we analyze each component systematically. Readers primarily interested in the final simplified equation may proceed to Eq. (S6).
  • Adiabatic Compression The momentum gain from plasma compression is:
@@
    $\dot{p}_{\rm adi}=-p(\nabla\cdot\vec{v}_{b}\,)/3\approx 1.03\times 10^{-6}\,{\rm GeV\,s^{-1}}\left(\frac{p}{1\mathrm{GeV}}\right)\left(\frac{r}{r_{g}}\right)^{-1}\left(\frac{v_{b}}{0.1}\right)\,$, (S2)
    where $r_{g}=3.1\times 10^{-4}$ pc is the gravitational radius of M87, and we approximate $-\nabla\cdot\vec{v}_{b}\approx v_{b}/r$.
  • Radiative Losses The total radiative loss encompasses four processes: $\dot{p}_{\rm rad}=\dot{p}_{\rm syn}+\dot{p}_{\rm brem}+\dot{p}_{\rm IC}+\dot{p}_{\rm C}$, which include synchrotron radiation ($\dot{p}_{\rm syn}$), bremsstrahlung ($\dot{p}_{\rm brem}$), inverse Compton scattering ($\dot{p}_{\rm IC}$), and Coulomb collisions ($\dot{p}_{\rm C}$). Their approximations are given as Regis and Ullio (2008); Chael et al. (2017):
[…]
    Here, we consider typical astrophysical conditions, including a magnetic field strength of $B\sim 1\,\rm G$, a background ion and thermal electron number density of $\bar{n}_{ion/e}\sim 1\,\rm cm^{-3}$, and a background radiation energy density of $\bar{u}_{\rm r}\sim 1\,\rm eV\,cm^{-3}$. Under these conditions, the synchrotron energy loss dominates, allowing us to neglect the latter three effects.
  • Spatial Diffusion Unlike the plasmas in GRMHD simulations, where the dynamics are dominated by much heavier ions with larger Larmor radii, the lighter electron-positron plasmas considered here are far less susceptible to diffusion. The spatial diffusion scale, defined as the distance over which charged particles lose most of their energy while diffusing, is given by Regis and Ullio (2008):
@@
    $\displaystyle d_{L}\equiv\sqrt{\frac{D_{xx}p}{\dot{p}}}\approx 4.18\times 10^{-4}\,r_{g}\,\left(\frac{B}{1\mathrm{G}}\right)^{-3/2}\ll r_{g}$ (S4)
    Here, the spatial diffusion coefficient follows from Bohm diffusion as $D_{xx}=p/(3eB)$, where energy losses are primarily due to synchrotron radiation. Compared to synchrotron losses, the effect of spatial diffusion is suppressed by $d_{L}/r_{g}$ and can thus be neglected.
  • Momentum Diffusion Momentum diffusion, caused by magnetic turbulence, leads to stochastic acceleration by Alfvén waves, contributing to momentum gain. This process is expressed as Dermer et al. (1996):
    $\displaystyle\dot{p}_{\rm acc}=\frac{m_{e}}{p^{2}}\frac{\partial}{\partial p}\left(p^{2}D_{pp}\right)\sim m_{e}\beta_{A}^{2}\zeta\,r_{L}^{\alpha-2}\lambda_{\rm max}^{1-\alpha}p^{\alpha-1}\approx 1.52\times 10^{-9}\,{\rm GeV\,s^{-1}}\beta_{A}^{2}\left(\frac{\zeta}{0.1}\right)\left(\frac{10^{4}\,r_{g}}{\lambda_{\rm max}}\right)\left(\frac{p}{\mathrm{GeV}}\right)\,.$ (S5)
    Here, the momentum diffusion coefficient is given by $D_{pp}\sim\beta_{A}^{2}\zeta r_{L}^{\alpha-2}\lambda_{\rm max}^{1-\alpha}p^{\alpha}$, where the magnetic turbulence follows a power-law spectrum with index $\alpha=2$. The parameter $\beta_{A}\approx 1$ represents the Alfvén velocity, $\zeta\equiv\delta B^{2}/B^{2}\sim 0.1$ quantifies the ratio of turbulent to average magnetic field strength, $r_{L}\equiv m_{e}/(eB)$ is the electron Larmor radius, and $\lambda_{\rm max}\sim 10^{4}r_{g}$ denotes the maximum wavelength of Alfvén waves. This contribution remains small compared to $\dot{p}_{\rm adi}$ and $\dot{p}_{\rm syn}$ and can therefore be neglected.
@@
Synthesizing the discussion above, we retain only the first three terms in Eq. (S1) and adopt the isotropic momentum approximation, simplifying Eq. (S1) to
@@
The evolution of the electron-positron distribution follows specific streamlines determined by the bulk velocity $\vec{v}_{b}$. Due to their high charge-to-mass ratios relative to heavier ions in the accretion flow, electrons and positrons closely follow magnetic field lines. Consequently, the streamline direction $\hat{v}_{b}(\vec{r}\,)$ aligns with the magnetic field direction, denoted as $\hat{B}(\vec{r}\,)$. The bulk velocity profiles considered in this study will be discussed in detail in Secs. I.3 and I.4.
Next, we consider the solution to Eq. (S6). The momentum evolution along a streamline can be parameterized as
@@
which can be directly integrated to obtain
$\displaystyle p-p_{i}(s_{i})=\int_{s_{i}}^{s_{\vec{r}}}{\rm d}s\,\frac{\dot{p}_{\rm adi}+\dot{p}_{\rm syn}}{v_{b}}.$ (S8)
Here, $s$ represents the distance traveled along the streamline, starting from an injection point $s_{i}$ with initial momentum $p_{i}$ due to DM annihilation, to a final point $s_{\vec{r}}$ at location $\vec{r}$. For a given momentum $p$ at $\vec{r}$, the corresponding injection momentum $p_{i}$ can be determined along the streamline.
The left-hand side of Eq. (S6) can be rewritten as ${\rm d}f_{e}/{\rm d}s$ along the streamline. The solution is then obtained by summing the contributions from all upstream injection points $s_{i}$ propagating to $\vec{r}$:
$\begin{split}f_{e}(\vec{r},p)&=\int^{s_{\vec{r}}}_{s_{b}}{\rm d}s_{i}\,\frac{q\,(s_{i},p_{i}(s_{i}))}{v_{b}(s_{i})}\exp\left[\int^{s_{i}}_{s_{\vec{r}}}{\rm d}{s}\,\frac{4\,\dot{p}_{\rm syn}(s,p_{i}({s}))}{p_{i}({s})\,v_{b}({s})}\right]\,\\ &=\int^{s_{\vec{r}}}_{s_{b}}{\rm d}s_{i}\frac{q\,(s_{i},p_{i}(s_{i}))}{v_{b}(s_{i})}\left(\frac{p_{i}(s_{i})}{p}\right)^{4}G(s_{\vec{r}})\,G^{-1}(s_{i}).\end{split}$ (S9)
Here, $s_{b}$ denotes the most distant injection point along the streamline, serving as a boundary condition. The function $G$ is a form factor that satisfies
@@
The factor $G(s_{\vec{r}})\,G^{-1}(s_{i})$ in Eq. (S9) accounts for the effects of flow compression and expansion. As shown in Eq. (S11), $G$ contains two distinct contributions: the first term arises from the convergence or divergence of $\hat{v}_{b}$, which is aligned with the magnetic field lines, while the second term captures the influence of bulk velocity variations on particle density. In the case of a spherically symmetric magnetic field and bulk velocity similar to free-fall, $G(s_{\vec{r}})\propto r^{-2}$, consistent with the results in Refs. Aloisio et al. (2004); Regis and Ullio (2008).
Notice that the above discussion is not in fully general relativistic setup. We will discuss the unceratinties arising from this treatment in the last section, which demonstrates that the approximation we employed is robust for our constraints.